OpenStudy (anonymous):

At time t1=12 s, a car is located at ⟨107,0,0⟩ m and has velocity ⟨5,0,0⟩ m/s. At time t2=16 s, what is the x component of the position of the car (in meters)?

OpenStudy (anonymous):

\[x2=(127,0,0)\]i think:\[\Delta x/\Delta t=v\]\[\Delta x=4(5,0,0)\]\[\Delta x=(20,0,0)\] x2=(127,0,0)

OpenStudy (anonymous):

Thanks !

OpenStudy (anonymous):

Question 5 Here are the positions at three different times for a bee in flight (a bee's top speed is about 7 m/s). time 5.4 s 5.9 s 6.4 s position ⟨−3.9,0,0⟩ m ⟨−1.4,0,0⟩ m ⟨0.6,0,0⟩ m (a) Compute the bee's average velocity between 5.4 s and 5.9 s. Be careful with signs. What was the x component of the bee's average velocity (in m/s) between 5.4 s and 5.9 s?

OpenStudy (anonymous):

@fazeli

OpenStudy (anonymous):

like up , we have:\[\Delta x /\Delta t=v\]\[\frac{ -1.4-(-3.9) }{ 5.9-5.4 }=5 m/s\]

OpenStudy (anonymous):

Thanks !

OpenStudy (anonymous):

finish?

OpenStudy (anonymous):

Compute the bee's average velocity between 5.4 s and 6.4 s. What was the x component of the bee's average velocity (in m/s) between 5.4 s and 6.4 s? Answer for Question 8

OpenStudy (anonymous):

Last one !

OpenStudy (anonymous):

like up:\[\frac{ 0.6-(-3.9) }{ 6.4-5.4 } =v=4.5 m/s\]